OpenStudy (anonymous):

Given quadrilateral ABCD ~ quadrilateral JKLM and AD = 14, JM = 6, and KL = 9, what is BC?

OpenStudy (er.mohd.amir):

sides are same ratio so AD/JM=BC/KL

OpenStudy (er.mohd.amir):

put all value and find BC

OpenStudy (anonymous):

divide?

OpenStudy (er.mohd.amir):

14/6=BC/9 14*9=6*BC BC=14*9/6 BC=21